OpenStudy (anonymous):

Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. -185 -77 -23 85

OpenStudy (anonymous):

3*3-2*5*2-3*3*3+-3*4

OpenStudy (anonymous):

27-50-81+81 = -23

OpenStudy (anonymous):

im assuming 3x3 and z4 is 3x^3 and z^4?

OpenStudy (anonymous):

sub the values of x, y and z into the equation (3)^3 - 2(5)^2 - 3(3)^3 + (-3)^4 = 27 - 2(25) -3(27) + 81 = 27 - 50 - 81 + 81 = -23